OpenStudy (anonymous):

Solve for x. 4 - 3x = 2x + 9 -13 -5 -1 1

OpenStudy (lgbasallote):

so where do you find it hard?

OpenStudy (anonymous):

4-3x=2x+9 -3x=2x+5 5x+5=0 5x=-5 x=-1 3rd One